Suppose Qd=-32*P+858, Qs=28*P+12 for home and Qd*=-32*P+172, Qs*=28*P+36 for foreign. What is the free trade volume to two decimal places?
Answer #1

Answer:

Under free trade, Exports of one country equals Import of the other.

For Home:

Qd = -32P + 858

Qs = 28P + 12

Imports = Qd - Qs = -32P + 858 - 28P - 12

Imports = - 60P + 846

For Foreign:

Qd = -32P + 172

Qs = 28P + 36

Exports = Qs* - Qd* = 28P + 36 – [-32P + 172] = 28P + 36 + 32P -172 = 60P -136

Equating Home Imports with Foreign Exports,

- 60P + 846 = 60P -136

120 P = 982

P = 982/120 = 8.18

Trade volume = 60P -136 = (60 x 8.18) - 136 = 490.8 - 136 = 354.8 = 355(aprrox.)

Thus trade volume is 355(approx.)